OpenStudy (anonymous):

Rhett decides to build a square room for his movie and music collection. If the area of the room is 9x2 − 6x + 1 square feet, what is the length of one side of the room? (3x + 1) feet (9x − 1) feet (3x − 1) feet (9x + 1) feet

OpenStudy (anonymous):

(3x-1) feet. Because the room is square the area=(length)*(length) So area=9x^2-6x+1=(length)^2 By factorising 9x^2-6x+1 you get (3x-1)(3x-1)=(3x-1)^2=(length)^2 Therefore length=(3x-1)